# Moments of characteristic polynomials and their derivatives for $SO(2N)$ and $USp(2N)$ and their application to one-level density in families of elliptic curve $L$-functions

### v1

@@ 2.1.3 Setting up the case of one pole away from zero
We shall consider the case where exactly one of the integrals is integrated around ±$\pm\alpha$, and the rest are integrated around zero. Let us first integrate $w_{1},\dots,w_{K-1}$ around zero and $w_{K}$ around $\alpha$, and call this term $\mathcal{T}^{SO(2N)}_{1}$. There will be $K$ other terms exactly equivalent to this term (i.e. identical up to relabelling of the integration variables) contributing to our final result, and $K$ terms equivalent to to integrating $w_{1},\dots,w_{K-1}$ around zero and $w_{K}$ around $-\alpha$. If we call the latter $\mathcal{T}^{SO(2N)}_{2}$, then the total contribution to $\mathcal{R}_{SO(2N)}(K,\alpha,\gamma)$ from the terms with one pole away from zero will be $K\mathcal{T}^{SO(2N)}_{1}+K\mathcal{T}^{SO(2N)}_{2}$. Let us find $\mathcal{T}^{SO(2N)}_{1}$ first, where we shall start by separating the $w_{K}$ variable from the products:
@@
×$\times\prod_{n=1}^{K}\left(\frac{\exp(Nw_{n})w_{n}^{3-2K}}{z\left(w_{n}+\gamma\right)\left(w_{n}-\alpha\right)\left(w_{n}+\alpha\right)}\right)\;\;\mathrm{d}w_{1}\cdots\mathrm{d}w_{K}$ (40)
@@
×$\times\prod_{n=1}^{K-1}\frac{\exp(Nw_{n})w_{n}^{3-2K}}{z(w_{n}+\gamma)(w_{n}-\alpha)(w_{n}+\alpha)}$
×$\times\left(\oint_{\alpha}\left[\prod_{j=1}^{K-1}z(w_{j}+w_{K})(w_{K}^{2}-w_{j}^{2})^{2}\right]\frac{\exp(Nw_{K})w_{K}^{3-2K}}{z(w_{K}+\gamma)(w_{K}-\alpha)(w_{K}+\alpha)}\mathrm{d}w_{K}\right)\mathrm{d}w_{1}\cdots\mathrm{d}w_{K-1}.$
@@
×$\times\oint_{0}\dots\oint_{0}\displaystyle\prod_{1\leq j<k\leq K-1}\left[\frac{1}{w_{j}+w_{k}}+\frac{1}{2N}+\mathcal{O}\left(N^{-2}\right)\right]\Delta\left(w_{1}^{2},\dots,w_{K-1}^{2}\right)^{2}$
×$\times\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\displaystyle\prod_{n=1}^{K-1}\left[\left(z(\alpha)+\frac{w_{n}}{N}z(\alpha)z(-\alpha)\right)\left(-\alpha^{2}\right)\frac{1}{z(\gamma)}\right.$
×$\times\left.\left(1-\frac{w_{n}}{N}z(-\gamma)\right)w_{n}^{3-2K}\mathrm{d}w_{n}\left(1+\mathcal{O}\left(N^{-2}\right)\right)\right]$
=\ = $(-1)^{K-1}N^{(K-1)(K-2)/2}C_{\alpha,\gamma,K}^{SO(2N)}(\pi\mathrm{i})\frac{\mathrm{e}^{N\alpha}}{z\left(\alpha+\gamma\right)}$
×$\times\oint_{0}\dots\oint_{0}\displaystyle\prod_{1\leq j<k\leq K-1}\left[\frac{1}{w_{j}+w_{k}}+\frac{1}{2N}\right]\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\Delta\left(w_{1}^{2},\dots,w_{K-1}^{2}\right)^{2}$
@@
×$\times\left[\oint\dots\oint\displaystyle\prod_{1\leq j<k\leq K-1}\left(\frac{1}{w_{j}+w_{k}}\right)\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\Delta\left(w_{1}^{2},\dots,w_{K-1}^{2}\right)^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}\right.$
@@
×$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}$
@@
×$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}$
@@
×$\times\left.\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}\right].$ (47)
Let
$\mathcal{M}^{SO(2N)}=\oint\dots\oint\displaystyle\prod_{1\leq j<l\leq K-1}\left(\frac{1}{w_{j}+w_{l}}\right)\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)$
×$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n},$ (48)
$\mathcal{J}^{SO(2N)}=\oint\dots\oint\left(\displaystyle\sum_{m=1}^{K-1}w_{m}\right)\displaystyle\prod_{1\leq j<l\leq K-1}\left(\frac{1}{w_{j}+w_{l}}\right)\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)$
×$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}.$ (49)
@@ 2.1.4 Evaluating $\mathcal{M}^{SO(2N)}$
×$\qquad\times w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K-1}^{1-K}\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K-1},$
where for the second line we have used the identical method to Section 2.1.2 (see the similarities with (2.1.2) - the difference being that we have just $K-1$ variables).
Continuing to follow the method from Section 2.1.2 we write $\mathcal{M}^{SO(2N)}$ as a determinant of gamma functions. Then, in the second line below, we write the gamma functions as factorials, and then reverse the order of the columns (acquiring an overall sign from the $(K-1)(K-2)/2$ column exchanges). We then multiply the new $j$th column by $(2j-2)!$:
@@
×$\qquad\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{ccccc}\frac{0!}{0!}&\cdots&\frac{(2K-8)!}{(2K-8)!}&\frac{(2K-6)!}{(2K-6)!}&\frac{(2K-4)!}{(2K-4)!}\\ \frac{0!}{(-1)!}&\cdots&\frac{(2K-8)!}{(2K-9)!}&\frac{(2K-6)!}{(2K-7)!}&\frac{(2K-4)!}{(2K-5)!}\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ \frac{0!}{(2-K)!}&\cdots&\frac{(2K-8)!}{(K-6)!}&\frac{(2K-6)!}{(K-4)!}&\frac{(2K-4)!}{(K-2)!}\end{array}\right).$
We now use a trick from [2] to turn this matrix into a Vandermonde determinant. We note that the first row above is all 1’s. The element in the $j$th position in row two is $22j-$ 2. Thus we can reduce the $j$th element in the second row to just $j$ by adding twice the first row to the second row (this does not change the determinant) and then pulling a factor of 2 out of the second row. Note that this procedure works equally well for the first entry in row two, which is zero.
@@
×$\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \frac{0!}{(-2)!}&\frac{2!}{0!}&\cdots&\frac{(2K-8)!}{(2K-10)!}&\frac{(2K-6)!}{(2K-8)!}&\frac{(2K-4)!}{(2K-6)!}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \frac{0!}{(2-K)!}&\frac{2!}{(4-K)!}&\cdots&\frac{(2K-8)!}{(K-6)!}&\frac{(2K-6)!}{(K-4)!}&\frac{(2K-4)!}{(K-2)!}\end{array}\right).$
We note that in the rest of the matrix the $j$th entry in the $m$th row is a polynomial in $j$ of order $m-1$ with leading coefficient $2^{m-1}$: for example, the $j$th entry in the third row is $(2j-2)(2j-3)$. Working one row at a time, all but the leading order term of the polynomial can be removed by adding multiples of rows higher up the matrix, and the power of 2 is pulled out of the determinant. This leaves us with
@@
×$\quad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&2^{K-2}&\cdots&(K-3)^{K-2}&(K-2)^{K-2}&(K-1)^{K-2}\end{array}\right)$
@@ 2.1.5 Evaluating $\mathcal{J}^{SO(2N)}$
×$\qquad\times w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K-1}^{1-K}\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K-1}.$
Using the usual method from Section 2.1.2 (noting that the integrand of (86) is symmetric as long as we retain the sum over $m$ intact) to pull all the factors containing $w_{j}$ into the $j$th row of the determinant, and also bringing the integration inside the determinant and evaluating the matrix elements as gamma functions, we see that each term in the sum over $m$ gives zero except for the $m=K-1$ term. This is because all terms except the $K-1$th raise the power on $w_{m}$, $m=1,\ldots,K-2$ by one, making the power equal to that on $w_{j+1}$. This results in two identical rows of gamma functions, and so the determinant is zero. For $m=K-1$, however, we get the following non-zero determinant, which has a jump of 2 in the argument of the gamma function between the second-to-last and last row instead of the usual change by 1:
@@
×$\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}\frac{0!}{0!}&\frac{2!}{2!}&\cdots&\frac{(2K-8)!}{(2K-8)!}&\frac{(2K-6)!}{(2K-6)!}&\frac{(2K-4)!}{(2K-4)!}\\ \frac{0!}{(-1)!}&\frac{2!}{1!}&\cdots&\frac{(2K-8)!}{(2K-9)!}&\frac{(2K-6)!}{(2K-7)!}&\frac{(2K-4)!}{(2K-5)!}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \frac{0!}{(3-K)!}&\frac{2!}{(5-K)!}&\cdots&\frac{(2K-8)!}{(K-5)!}&\frac{(2K-6)!}{(K-3)!}&\frac{(2K-4)!}{(K-1)!}\\ \frac{0!}{(1-K)!}&\frac{2!}{(3-K)!}&\cdots&\frac{(2K-8)!}{(K-7)!}&\frac{(2K-6)!}{(K-5)!}&\frac{(2K-4)!}{(K-3)!}\end{array}\right).$
As in the previous section, we note that for the first $K-2$ rows the matrix element in the $j$th column and $m$th row is a polynomial in $j$ of order $m-1$. For example, the $j$th element in the 2nd row is just $22j-$ 2 and the $j$th element in the $(K-2)$th row is $(2j-2)(2j-3)\cdots(2j-(K-2))$. Working sequentially from the top, each polynomial can be reduced to just its highest order term, $(2j)^{m-1}$ by adding multiples of previous rows. The final row, the only one that differs from (73), contains polynomials of order $K-1$: $(2j-2)(2j-3)\ldots(2j-K)=(2j)^{K-1}-(2j)^{K-2}(2+3+\cdots+K)+\mathcal{O}(j^{K-3})$. So, splitting the determinant into two determinants, each one containing just one of the two surviving terms from the bottom row, we have
@@
×$\qquad\times\left[2\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1^{K-3}&2^{K-3}&\cdots&(K-3)^{K-3}&(K-2)^{K-3}&(K-1)^{K-3}\\ 1^{K-1}&2^{K-1}&\cdots&(K-3)^{K-1}&(K-2)^{K-1}&(K-1)^{K-1}\end{array}\right)\right.$ (106)
@@
×$\qquad\;\;\times\left.\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1^{K-3}&2^{K-3}&\cdots&(K-3)^{K-3}&(K-2)^{K-3}&(K-1)^{K-3}\\ 1^{K-2}&2^{K-2}&\cdots&(K-3)^{K-2}&(K-2)^{K-2}&(K-1)^{K-2}\end{array}\right)\right].$ (112)
[…]
@@
×$\qquad\times\frac{(K-1)(K-2)}{2}\Delta(1,2,3,\ldots,K-1)$
@@
By identical methods it can be shown that the third integral in (2.1.3) is
@@
×$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}$

### v2

@@ 2.1.3 Setting up the case of one pole away from zero
We shall consider the case where exactly one of the integrals is integrated around $\pm\alpha$, and the rest are integrated around zero. Let us first integrate $w_{1},\dots,w_{K-1}$ around zero and $w_{K}$ around $\alpha$, and call this term $\mathcal{T}^{SO(2N)}_{1}$. There will be $K$ other terms exactly equivalent to this term (i.e. identical up to relabelling of the integration variables) contributing to our final result, and $K$ terms equivalent to to integrating $w_{1},\dots,w_{K-1}$ around zero and $w_{K}$ around $\alpha$. If we call the latter $\mathcal{T}^{SO(2N)}_{2}$, then the total contribution to $\mathcal{R}_{SO(2N)}(K,\alpha,\gamma)$ from the terms with one pole away from zero will be $K\mathcal{T}^{SO(2N)}_{1}+K\mathcal{T}^{SO(2N)}_{2}$. Let us find $\mathcal{T}^{SO(2N)}_{1}$ first, where we shall start by separating the $w_{K}$ variable from the products:
@@
$\times\prod_{n=1}^{K}\left(\frac{\exp(Nw_{n})w_{n}^{3-2K}}{z\left(w_{n}+\gamma\right)\left(w_{n}-\alpha\right)\left(w_{n}+\alpha\right)}\right)\;\;\mathrm{d}w_{1}\cdots\mathrm{d}w_{K}$ (40)
@@
$\times\prod_{n=1}^{K-1}\frac{\exp(Nw_{n})w_{n}^{3-2K}}{z(w_{n}+\gamma)(w_{n}-\alpha)(w_{n}+\alpha)}$
$\times\left(\oint_{\alpha}\left[\prod_{j=1}^{K-1}z(w_{j}+w_{K})(w_{K}^{2}-w_{j}^{2})^{2}\right]\frac{\exp(Nw_{K})w_{K}^{3-2K}}{z(w_{K}+\gamma)(w_{K}-\alpha)(w_{K}+\alpha)}\mathrm{d}w_{K}\right)\mathrm{d}w_{1}\cdots\mathrm{d}w_{K-1}.$
@@
$\times\oint_{0}\dots\oint_{0}\displaystyle\prod_{1\leq j<k\leq K-1}\left[\frac{1}{w_{j}+w_{k}}+\frac{1}{2N}+\mathcal{O}\left(N^{-2}\right)\right]\Delta\left(w_{1}^{2},\dots,w_{K-1}^{2}\right)^{2}$
$\times\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\displaystyle\prod_{n=1}^{K-1}\left[\left(z(\alpha)+\frac{w_{n}}{N}z(\alpha)z(-\alpha)\right)\left(-\alpha^{2}\right)\frac{1}{z(\gamma)}\right$.
$\times\left.\left(1-\frac{w_{n}}{N}z(-\gamma)\right)w_{n}^{3-2K}\mathrm{d}w_{n}\left(1+\mathcal{O}\left(N^{-2}\right)\right)\right]$
$=\$$(-1)^{K-1}N^{(K-1)(K-2)/2}C_{\alpha,\gamma,K}^{SO(2N)}(\pi\mathrm{i})\frac{\mathrm{e}^{N\alpha}}{z\left(\alpha+\gamma\right)}$
$\times\oint_{0}\dots\oint_{0}\displaystyle\prod_{1\leq j<k\leq K-1}\left[\frac{1}{w_{j}+w_{k}}+\frac{1}{2N}\right]\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\Delta\left(w_{1}^{2},\dots,w_{K-1}^{2}\right)^{2}$
@@
$\times\left[\oint\dots\oint\displaystyle\prod_{1\leq j<k\leq K-1}\left(\frac{1}{w_{j}+w_{k}}\right)\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\Delta\left(w_{1}^{2},\dots,w_{K-1}^{2}\right)^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}\right$.
@@
$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}$
@@
$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}$
@@
$\times\left.\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}\right].$ (47)
Let us label the integrals in (2.1.3):
$\mathcal{M}^{SO(2N)}=\oint\dots\oint\displaystyle\prod_{1\leq j<k\leq K-1}\left(\frac{1}{w_{j}+w_{k}}\right)\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)$
$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n},$ (48)
$\mathcal{J}^{SO(2N)}=\oint\dots\oint\left(\displaystyle\sum_{m=1}^{K-1}w_{m}\right)\displaystyle\prod_{1\leq j<k\leq K-1}\left(\frac{1}{w_{j}+w_{k}}\right)\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)$
$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}.$ (49)
Note that the final integral in (2.1.3) is just $(K-2)\mathcal{J}^{SO(2N)}$, and in fact we will see in (114) that $\mathcal{J}^{SO(2N)}$ and $\mathcal{M}^{SO(2N)}$ are related by $\mathcal{J}^{SO(2N)}=\frac{(K-1)(K-2)}{2}\mathcal{M}^{SO(2N)}$.
@@ 2.1.4 Evaluating $\mathcal{M}^{SO(2N)}$
$\qquad\times w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K-1}^{1-K}\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K-1},$
where for the second line we have used the identical method to Section 2.1.2 (see the similarities with (2.1.2) - the difference being that we have just K1K-1 variables).
Continuing to follow the method from Section 2.1.2 we write $\mathcal{M}^{SO(2N)}$ as a determinant of gamma functions. Then, in the second line below, we write the gamma functions as factorials, and then reverse the order of the columns (acquiring an overall sign from the (K1)(K2)/2(K-1)(K-2)/2 column exchanges). We then multiply the new $j$th column by (2j2)!(2j-2)!:
@@
$\qquad\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{ccccc}\frac{0!}{0!}&\cdots&\frac{(2K-8)!}{(2K-8)!}&\frac{(2K-6)!}{(2K-6)!}&\frac{(2K-4)!}{(2K-4)!}\\ \frac{0!}{(-1)!}&\cdots&\frac{(2K-8)!}{(2K-9)!}&\frac{(2K-6)!}{(2K-7)!}&\frac{(2K-4)!}{(2K-5)!}\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ \frac{0!}{(2-K)!}&\cdots&\frac{(2K-8)!}{(K-6)!}&\frac{(2K-6)!}{(K-4)!}&\frac{(2K-4)!}{(K-2)!}\end{array}\right).$
We now use a trick from [2] to turn this matrix into a Vandermonde determinant. We note that the first row above is all 1’s. The element in the $j$th position in row two is 2j22j-2. Thus we can reduce the $j$th element in the second row to just $j$ by adding twice the first row to the second row (this does not change the determinant) and then pulling a factor of 2 out of the second row. Note that this procedure works equally well for the first entry in row two, which is zero.
@@
$\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \frac{0!}{(-2)!}&\frac{2!}{0!}&\cdots&\frac{(2K-8)!}{(2K-10)!}&\frac{(2K-6)!}{(2K-8)!}&\frac{(2K-4)!}{(2K-6)!}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \frac{0!}{(2-K)!}&\frac{2!}{(4-K)!}&\cdots&\frac{(2K-8)!}{(K-6)!}&\frac{(2K-6)!}{(K-4)!}&\frac{(2K-4)!}{(K-2)!}\end{array}\right).$
We note that in the rest of the matrix the $j$th entry in the $m$th row is a polynomial in $j$ of order m1m-1 with leading coefficient $2^{m-1}$: for example, the $j$th entry in the third row is (2j2)(2j3)(2j-2)(2j-3). Working one row at a time, all but the leading order term of the polynomial can be removed by adding multiples of rows higher up the matrix, and the power of 2 is pulled out of the determinant. This leaves us with
@@
$\quad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&2^{K-2}&\cdots&(K-3)^{K-2}&(K-2)^{K-2}&(K-1)^{K-2}\end{array}\right)$
@@ 2.1.5 Evaluating $\mathcal{J}^{SO(2N)}$
$\qquad\times w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K-1}^{1-K}\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\mathrm{d}w_{1}\dots\mathrm{d}w_{K-1}.$
Using the usual method from Section 2.1.2 (noting that the integrand of (86) is symmetric as long as we retain the sum over $m$ intact) to pull all the factors containing $w_{j}$ into the $j$th row of the determinant, and also bringing the integration inside the determinant and evaluating the matrix elements as gamma functions, we see that each term in the sum over $m$ gives zero except for the m=K1m=K-1 term. This is because all terms except the K1K-1th raise the power on $w_{m}$, $m=1,\ldots,K-2$ by one, making the power equal to that on $w_{j+1}$. This results in two identical rows of gamma functions, and so the determinant is zero. For m=K1m=K-1, however, we get the following non-zero determinant, which has a jump of 2 in the argument of the gamma function between the second-to-last and last row instead of the usual change by 1:
@@
$\qquad\times\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}\frac{0!}{0!}&\frac{2!}{2!}&\cdots&\frac{(2K-8)!}{(2K-8)!}&\frac{(2K-6)!}{(2K-6)!}&\frac{(2K-4)!}{(2K-4)!}\\ \frac{0!}{(-1)!}&\frac{2!}{1!}&\cdots&\frac{(2K-8)!}{(2K-9)!}&\frac{(2K-6)!}{(2K-7)!}&\frac{(2K-4)!}{(2K-5)!}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \frac{0!}{(3-K)!}&\frac{2!}{(5-K)!}&\cdots&\frac{(2K-8)!}{(K-5)!}&\frac{(2K-6)!}{(K-3)!}&\frac{(2K-4)!}{(K-1)!}\\ \frac{0!}{(1-K)!}&\frac{2!}{(3-K)!}&\cdots&\frac{(2K-8)!}{(K-7)!}&\frac{(2K-6)!}{(K-5)!}&\frac{(2K-4)!}{(K-3)!}\end{array}\right).$
As in the previous section, we note that for the first K2K-2 rows the matrix element in the $j$th column and $m$th row is a polynomial in $j$ of order m1m-1. For example, the $j$th element in the 2nd row is just 2j22j-2 and the $j$th element in the (K2)(K-2)th row is $(2j-2)(2j-3)\cdots(2j-(K-2))$. Working sequentially from the top, each polynomial can be reduced to just its highest order term, $(2j)^{m-1}$ by adding multiples of previous rows. The final row, the only one that differs from (73), contains polynomials of order K1K-1: $(2j-2)(2j-3)\ldots(2j-K)=(2j)^{K-1}-(2j)^{K-2}(2+3+\cdots+K)+\mathcal{O}(j^{K-3})$. So, splitting the determinant into two determinants, each one containing just one of the two surviving terms from the bottom row, we have
@@
$\qquad\times\left[2\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1^{K-3}&2^{K-3}&\cdots&(K-3)^{K-3}&(K-2)^{K-3}&(K-1)^{K-3}\\ 1^{K-1}&2^{K-1}&\cdots&(K-3)^{K-1}&(K-2)^{K-1}&(K-1)^{K-1}\end{array}\right)\right$. (106)
@@
$\qquad\;\;\times\left.\det_{(K-1)\times(K-1)}\left(\begin{array}[]{cccccc}1&1&\cdots&1&1&1\\ 1&2&\cdots&(K-3)&(K-2)&(K-1)\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1^{K-3}&2^{K-3}&\cdots&(K-3)^{K-3}&(K-2)^{K-3}&(K-1)^{K-3}\\ 1^{K-2}&2^{K-2}&\cdots&(K-3)^{K-2}&(K-2)^{K-2}&(K-1)^{K-2}\end{array}\right)\right].$ (112)
The second determinant above is the Vandermonde $\Delta(1,2,3,\ldots K-1)=\prod_{j=1}^{K-1}(j-1)!$, while the first determinant is what is called in [26] a “Vandemondian” (the notation for this matrix in that paper is $V_{K-1\;\;1}$). Theorem I of [26] tells us that
@@
$\qquad\times\frac{(K-1)(K-2)}{2}\Delta(1,2,3,\ldots,K-1)$
@@
We have already seen that the third integral in (2.1.3) is
@@
$\times\Delta(w_{1}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{n=1}^{K-1}w_{n}^{3-2K}\mathrm{d}w_{n}$